Calculate the pH of each of the following solutions of a strong acid in water: a. 1.0 x 10-11 M HCI b. 5.0 M HCI
Answer #1

Here strong acid is given so we will consider complete dissociation.

pH is given as negative logarithm of hydrogen ion concentration i.e; pH = - log[H+]

Here concentrations are given in terms of molarity.

# by mole concept we can say, [HCl] =[ H+ ​​​​​​] ​​i.e; concentration of acid is equal to concentration of H+ given by the acid. So,

[HCl] = [H+] = 1-0 x 10-11M =-log (1.OND) b) 5.0 M HQ